OpenStudy (anonymous):

find the coordinates of missing endpoint given that N is the midpoint of MO. M(5,8) N(4,3)

OpenStudy (aroub):

(3+8/2)(4+5/2) =(11/2) (9/2)

OpenStudy (amistre64):

N(4 , 3) -M(5 , 8) --------- (-1,-5) +N( 4 , 3 ) ---------- O ( 3 ,-2 )

OpenStudy (anonymous):

(5.5,4.5)

OpenStudy (anonymous):

aroub?

OpenStudy (amistre64):

5(-1) = 4(-1) = 3 M N O 8(-5) = 3(-5) = -2

OpenStudy (aroub):

umm yes shericar :)

myininaya (myininaya):

\[(\frac{5+x}{2}=4,\frac{8+y}{2}=3)\] (5+x=8,8+y=6) (x=3,y=-2)

OpenStudy (amistre64):

-1 -1 <....5.....4......3...> M N O <....8.....3.....-2...> -5 -5

myininaya (myininaya):

amistre gj